OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = x + 3.

hero (hero):

Hint: The slope of the perpendicular line is the negative reciprocal of the given line.

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7 i got b. is it right?

hero (hero):

Weird. None of those answer choices include the negative reciprocal of the slope of the given line. The product of the slopes of two lines perpendicular to each other is -1.

OpenStudy (anonymous):

please double check your question